(1) any two groups of order m are isomorphic where m€ N
(2)if A and B are two sets such that A U B = phi , then A intersection B = phi
True/false. Justify.
1
Expert's answer
2020-03-30T11:07:18-0400

1. False.

Consider the group "D_4 =<r,s \mid r^4=s^2=1,rs=sr^{-1} >".

and "\Z_4=\{0,1,2,3,4,5,6,7 \}"

Where "D_4" is dihedral group of order 8 and "\Z_8" is a additive group of integer modulo 8.

Both "D_4 \ and \ \Z_8" have order 8 but "D_4" is not isomorphic to "\Z_8"

because "D_4" is non abelian and "\Z_8" is abelian.


2.True.

Since "A \cup B=\{ x:x\in A \ or \ x\in B \}=\phi"

Therefore,"A=B=\phi" .

Hence,"A\cap B=\phi" .